A substance has a density of 1.23 g/ml and a mass of 86.4 grams. What is the volume of the substance in ml?


1
Expert's answer
2021-03-22T06:21:54-0400

Density,d= 1.23g/ml

Mass,m= 86.4g

Volume,V= ?

d= m/V

V= m/d= 86.4/1.23= 70.24ml